OpenStudy (anonymous):

What are the x- and y- intercepts of the following equation: 8x - 4y = 12

OpenStudy (anonymous):

x-intercept, you plug y = 0. y-intercept, you plug x = 0

OpenStudy (anonymous):

x-intercept: 8x - 4(0) = 12 8x = 12 x = 12/8 x = 3/2 (3/2, 0) y-intercept: 8(0) - 4y = 12 -4y = 12 y = -12/4 y = -3 (0, -3)

OpenStudy (anonymous):

I hope you know the concept.